• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Abdurahim

    Log entry time 08:23:08 on April21,2010

    Entry number 318594

    keyword=Owl shift summary

    April 21, 2010


    SL: Abdurahim Rakhman


    TO: Abdurahim Rakhman

    3rd: Hisham Albataineh


    Swing Shift Summary:


    ++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++


    Plan: Compton running and production running.


    00:34 Stopped Compton run and increased GSO HV to -1700 V. Signal to noise ~ 1.4.

    02:20 There is no beam

    02:25 Called MCC and learned that they are having an RF problem and been told they will resume very soon.

    02:29 Beam is back

    03:52 Start doing Compton photon beam polarization scan. Set the QWP to an angle turning the cavity lock ON and OFF. Stayed at -1700 volts at PMT HV.

    05:30 Observed "BLumi Signals Around" distribution in online analysis gui is having two peaks.

    05:35 Compton PMT HV tripped. Turning the crate OFF and ON fixed the problem.

    05:50 By looking at the BPM plot dithering may not be running. Turned off and on it according to the instruction sheet. Seems still not working.

    06:30 Called Bob on dithering problem. Asked to make a halog entry of what we did. Regarding the BLumi problem Bob said we don't use it for lead so we don't have to worry about it.

    06:54 Compton PMT HV trip. Resetting it fixed the problem.

    07:50 Dithering can be seen in BPM strip chart. Error massage in ROC23 terminal has disappeared.


    There was no major problem during the whole shift except for having couple of high voltage trips. During the shift we collected total 1.3 C charge.








    Parity Run Summary:







































     PREX Run #   Beam Current (uA)   # of Good Multiplets   Cummulative # of Good Multiplets   Charge (mC)   Cummulative Charge (mC)   Comments   IHWP 
    2996 50 82315 88246 150 161 Production OUT
    2997 50 85556 173802 157 318 Production OUT
    2998 50 82390 256192 153 471 Production OUT
    2999 50 71306 327498 133 604 Production OUT
    3000 50 79699 407197 146 750 Production OUT
    3001 50 84468 491665 154 904 Production OUT
    3002 50 85498 577163 156 1060 Production OUT
    3003 50 68089 645252 127 1187 Production OUT
    3004 50 82682 727934 153 1340 Production OUT
    3005 50 N/A N/A N/A N/A Production OUT





    Compton Run Summary:

































     Run #   Start Time   Comments   IHWP 
    22577 00:34 Production (-1.7 kV) OUT
    22578 01:28 Production (-1.7 kV) OUT
    22579 02:21 junk (no beam) OUT
    22580 02:29 Production (-1.7 kV) OUT
    22581 03:12 Production (-1.7 kV) OUT
    22582 03:52 QWP#1 at 309° (-1.7 kV) OUT
    22583 04:57 QWP#1 at 306.5° (-1.7 kV) OUT
    22584 06:15 QWP#1 at 311.5° (-1.7 kV) OUT
    22585 07:21 QWP#1 at 316.5° (-1.7 kV) OUT
    22586 08:20 QWP#1 at 321.5° (-1.7 kV) OUT